Lupo Corporation uses a job-order costing system with a single plantwide predetermined overhead rate based on machine-hours.thanks
Answer #1

Total Overhead rate = Variable rate + Fixed rate

= 2 + (581,400/32,300)

= 20 per machine hour

Total Cost for the job = Direct Materials + Direct Labor + Overhead

= 630 + 1260 + (40*20) = 2690

Estimated selling price = 2690 + 40% = 3766

Selling price per unit = 3766/10

= 376.60